OpenStudy (anonymous):

5(sqrt m)=3(sqrt m+9)

OpenStudy (anonymous):

\[5\sqrt{m}=3\sqrt{m+9}\]

OpenStudy (lgbasallote):

\[5\sqrt m = 3\sqrt{m+9}\] square both sides \[25m = 9(m+9)\] disribute \[25m = 9m + 81\] subtract 9 from both sides \[25m - 9m = 81\] \[16m = 81\] divide both sides by 16 \[m = \frac{81}{16}\] does that help?

OpenStudy (anonymous):

oops... this is what I meant \[5\sqrt{m}=\sqrt{3m+9}\]

OpenStudy (ganpat):

ha ha, now follow the same.. but putting 3 inside the sq rt sign..:D

OpenStudy (lgbasallote):

so what's your answer to this one @a.galvan2 ?

OpenStudy (anonymous):

m=81/16

OpenStudy (lgbasallote):

how did you get that?

OpenStudy (ganpat):

 m poor at writing...